# A Computationally Efficient Maximum A Posteriori Sequence Estimation via Stein Variational Inference

### v3

@@ VI Conclusion
State estimation in autonomous systems is a significant challenge, often involving complex scenarios and multimodal distributions in real-world settings. Our estimation methods are designed to handle these multimodal distributions efficiently. In this article, we introduced a MAP sequence estimation method that effectively manages multimodality while reducing computational and memory requirements. The approach is grounded in a sequential variational inference framework that accounts for dependencies between transition states over time in dynamic system models. Building on this foundation, we addressed MAP sequence estimation through a deterministic gradient flow of particles, embedded within a Viterbi-style dynamic programming framework. To validate our approach, we rigorously evaluated the proposed method in challenging multimodal scenarios, including nonlinear dynamics with ambiguous measurements, pose estimation under unknown data associations, and real-world range-only (wireless) localization under temporary unobservability. Our proposed Stein-MAP-Seq consistently demonstrated robustness across these diverse scenarios.

### v4

@@ VI Conclusion
State estimation in autonomous systems remains a fundamental challenge, particularly in real-world settings characterized by complex dynamics and highly multimodal uncertainty. In this work, we introduced a computationally efficient MAP sequence estimation framework that explicitly addresses multimodal posterior distributions while reducing computational and storage requirements. The proposed approach is grounded in a sequential variational inference formulation that explicitly accounts for temporal dependencies induced by system dynamics.
Building on this formulation, we addressed MAP sequence estimation through a parallelizable particle-based gradient flow embedded within a Viterbi-style dynamic programming framework. This combination enables efficient mode concentration and globally consistent trajectory selection over long horizons. We rigorously evaluated the proposed method in a range of challenging multimodal scenarios, including nonlinear systems with ambiguous measurements, pose estimation under unknown data association, range-only localization under temporary unobservability, and high-dimensional manipulator estimation with kinematic redundancy.
Across all scenarios, Stein-MAP-Seq consistently demonstrated strong robustness and scalability. Moreover, beyond empirical accuracy, this work provides practical design insights for MAP sequence estimation in multimodal settings. Our results indicate that MAP-Seq estimation benefits primarily from coherent mode concentration and temporal consistency, rather than accurate approximation of the full posterior distribution. Accordingly, Stein-MAP-Seq achieves reliable MAP trajectory recovery using a small number of finite states, with performance governed more by kernel-induced mode structure than particle count alone. Moreover, the proposed method complements batch MAP optimization by providing robust global initialization under strongly multimodal and nonconvex objectives, highlighting its utility both as a standalone estimator and as a front-end for existing optimization-based frameworks.
@@
-F Per-Step Runtime and Scalability Analysis
We examine the computational efficiency of (IV.3) in Stein-MAP-Seq by measuring the average per-step runtime on a GPU as a function of the number of particles $N_{s}$ and the state dimension $n_{x}$. For consistency with all experimental scenarios, a fixed number of SVGD update iterations per time step is used ($100$ iterations). This analysis illustrates how the computational cost grows as the number of particles and the state dimension increase, thereby providing practical evidence that Stein-MAP-Seq remains computationally viable in high-dimensional estimation problems.
(a) Effect of Particle Count on Per-Step Runtime
(b) Effect of State Dimension on Per-Step Runtime
Figure 14: Per-step computational cost of Stein-MAP-Seq, averaged over Monte Carlo runs, measured using a fixed number of SVGD update iterations per time step (100 iterations). The results illustrate stable runtime behavior across varying estimation problem sizes.